### v2

# A Gentle Wakeup Call Symmetry Breaking with Less Collision Cost

Abstract
The wakeup problem addresses the fundamental challenge of symmetry breaking. Initially, $n$ devices share a time-slotted multiple access channel, which models wireless communication. A transmission succeeds if exactly one device sends in a slot; if two or more transmit, a collision occurs and none succeed. The goal is to achieve a single successful transmission efficiently.
Prior work on wakeup primarily analyzes latency—the number of slots until the first success. However, in many modern systems, each collision incurs a nontrivial delay, $\mathcal{C}$, which prior analyses neglect. Consequently, although existing algorithms achieve polylogarithmic-in-$n$ latency, they still suffer a delay of $\Omega(\mathcal{C})$ due to collisions.
Here, we design and analyze a randomized wakeup algorithm, Aim-High. When $\mathcal{C}$ is sufficiently large with respect to $n$, Aim-High has expected latency and expected total cost of collisions that are nearly $O(\sqrt{\mathcal{C}})$; otherwise, both quantities are $O({\texttt{poly}({\log n})})$. Finally, for a well-studied class of algorithms, we establish a trade-off between latency and expected total cost of collisions.
1 Introduction
In the wakeup problem, devices are either active (awake) or inactive (asleep). Each active device has a packet to send on a shared channel, while inactive devices are limited to listening. The first successful transmission serves as a “wakeup call”, awakening all inactive devices and resolving the problem [2].
Prior work on this problem has focused on the time until the first success, which is typically referred to as the latency. Minimizing this metric is challenging, since in any unit of time, the shared channel can only accommodate a single sender; if two or more devices send simultaneously, the result is a collision where none of the corresponding packets are successfully sent. Sending on the channel is performed in a distributed fashion; that is, a priori there is no central scheduler or coordinator.
Under the standard cost model for wakeup, each collision wastes a single slot. However, in many settings, a collision can consume much more than a single slot. Therefore, collisions may significantly impact the performance of modern systems.
In this work, we address both latency and the cost of collisions required to solve the wakeup problem. While existing algorithms give strong latency guarantees, they result in a large number of collisions (see Section 3.1), and addressing these two metrics in tandem requires new ideas.
Why Should We Care About Collisions? The delay incurred by a collision can vary across different systems. As an example, consider WiFi (IEEE 802.11) networks; here, both sending a packet and the delay from a collision consumes time, say $\mathcal{C}$, proportional to the packet size [1]. In this setting, wakeup incurs an unavoidable delay of $\mathcal{C}$, since a single success is required. Yet, reducing the delay caused by collisions is still important. If this delay is, say, $X$, then throughput is $\mathcal{C}/(\mathcal{C}+X)$. So, if $X=\Theta(\mathcal{C})$, then throughput is sub-optimal by a constant factor. However, if say $X=\Theta(\sqrt{\mathcal{C}})$, then throughput is $\mathcal{C}/(\mathcal{C}+\Theta(\sqrt{\mathcal{C}}))\geq 1-\Theta(1/\sqrt{\mathcal{C}})$; that is, throughput approaches the optimal value of $1$ as $\mathcal{C}$ grows large. [^1] Empty slots will also degrade throughput, but in WiFi, these slots are negligible compared to the time of sending a packet or a collision. This illustrates how minimizing collisions improves performance.
More generally, over long-range links, where the propagation delays are large, a collision means devices can wait a long round-trip time before realizing the transmission has failed. Also, in intermittently-connected mobile wireless networks [21], failed communication due to a collision may result in the sender having to wait until the intended receiver is again within transmission range. Finally, in environments where there is some fluctuation in the delay (e.g., arising from fluctuating channel conditions or periods of heavy network usage), $\mathcal{C}$ can be interpreted as an upper bound on this delay that holds under most operating conditions.
1.1 Model and Notation
[…]
Communication. Time is divided into disjoint slots, each of which can accommodate a packet transmission. Communication occurs on a multiple access channel, which is defined as follows. In any slot, a packet can either attempt to transmit or listen to the channel. If no packet is transmitted in a given slot, it is empty. If only one packet is transmitted, the packet succeeds; we often refer to this as a success. However, if multiple packets transmit simultaneously, they all fail—this is a collision—and they may try again later. There is no scheduler or central authority a priori.
No Collision Detection. There is no mechanism to distinguish between an empty slot and a collision—that is, no collision detection (no-CD). In many practical systems, collision detection (CD) is unavailable or unreliable. For example, in WiFi networks, received signal strength may offer a weak form of CD, but it is error-prone. Consequently, many prior results address the challenging no-CD setting (see the survey [2] and related discussion in Section 2).
We emphasize that CD and the cost of collisions are orthogonal concerns: collisions incur a performance penalty regardless of whether they are detectable. For example, in WiFi, even if devices cannot sense an ongoing collision, the channel remains unavailable for its duration.
Metrics. We care about latency, which refers to the time between the first packet injection and the first success. Additionally, each collision incurs a known delay of $\mathcal{C}$ $\geq 4$. [^2] The value $4$ is chosen to simplify the analysis. We also note that only $\mathcal{C}=\omega(1)$ is interesting; otherwise, an optimal algorithm (such as sawtooth backoff [13, 10, 2]) for the standard model will be asymptotically optimal. To capture the worst-case performance, $\mathcal{C}$ is chosen by the adversary with knowledge of $n$, subject to $\mathcal{C}=O({\texttt{poly}({n})})$, where the degree is of the polynomial is unknown. Note that this reveals no information about $n$, whose value is unknown to the packets a priori.
Given any wakeup algorithm, the two pertinent costs are: (i) the total cost of collisions, which is the number of collisions incurred over the execution of the algorithm multiplied by $\mathcal{C}$, and (ii) the latency. We report on both of these metrics separately, with the aim of minimizing their maximum.
Notation. An event holds with high probability (w.h.p.) in $x$ if it occurs with probability at least $1-O(1/{\texttt{poly}({x})})$, where we can make the constant degree of the polynomial arbitrarily large. We use $\log^{c}y$ to denote $(\log y)^{c}$.
1.2 Our Results
The performance of our algorithm, Aim-High (AH), is parameterized by $\epsilon$, which is a small positive constant set by the user prior to execution.
Theorem 1.1
AH solves the wakeup problem with the following costs:
  1. 1.
    If $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, then the expected latency is $O(\mathcal{C}^{\frac{1}{2}+2\epsilon})$ and the expected total cost of collisions is $O(\mathcal{C}^{\frac{1}{2}+\epsilon})$.
  2. 2.
    Else, the expected latency is $O(\log^{4+\frac{1}{2\epsilon}}n)$ and the expected total cost of collisions is $O(\log^{4+\frac{3}{2\epsilon})}n)$.
By comparison, existing approaches to wakeup incur a collision cost that is $\Omega(\mathcal{C})$ (see Section 3.1). In contrast, Theorem 1.1 implies that if $\mathcal{C}$ is “large”, AH offers performance that is sublinear in $\mathcal{C}$. Else, latency and expected total cost of collisions are polylogarithmic-in-$n$, which echoes prior wakeup results (see Section 2).
Many upper bounds are achieved by “fair” algorithms. A fair algorithm is defined by a sequence $p_{1}$, $p_{2}$, $\dots$, where $p_{i}$ denotes the sending probability used by every packet in the $i$-th slot. In Section 5, we establish the following trade-off between latency and expected total cost of collisions.
Theorem 1.2
Let $\alpha$ and $\beta$ be constants such that $1/(2\epsilon)\geq\alpha\geq 0$ and $\beta\geq 0$. Any fair algorithm for the wakeup problem that has $O(\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n)$ latency has an expected total cost of collisions $\Omega(\mathcal{C}^{(1/2)-\alpha\epsilon}/\log^{\beta}n)$.
2 Related Work
Wakeup, or leader election, addresses the fundamental problem of symmetry breaking. Early work by Willard [20] establishes a $\lg\lg n+O(1)$ upper bound with a matching lower bound for fair algorithms under collision detection (CD).
In contrast, without CD, a common theme is that the latency increases to logarithmic or polylogarithmic in $n$. For starters, without collision detection and when $n$ is known, Kushilevitz and Mansour [16, 15] prove a lower bound of $\Omega(\log n)$ expected latency. Newport [18, 19], in work that unifies many prior lower-bound results, shows $\Omega(\log n)$ expected latency and $\Omega(\log^{2}n)$ latency w.h.p. for general randomized algorithms. Complementing these lower bounds, Bar-Yehuda et al. [3] achieve a matching upper bound of $O(\log n)$ expected latency and w.h.p. $O(\log^{2}n)$ latency.
Farach-Colton et al. [7] obtain $\Omega(\log n\log(1/\varepsilon))$ latency for algorithms without CD, succeeding with probability at least $1-\varepsilon$. A complementary by Jurdziński and Stachowiak [14] demonstrates an algorithm that succeeds with probability at least $1-\varepsilon$ with $O(\log n\log(1/\varepsilon))$ latency, without CD, but assuming unique labels or knowledge of $n$.
With multiple channels and collision detection, Fineman et al. [9] give an efficient contention-resolution algorithm. In the SINR model [17], Fineman et al. [8] achieve logarithmic latency and establish matching lower bounds. Chang et al. [6] examine energy efficiency under varying feedback models, while Gilbert et al. [11] study the effect of system-size prediction, deriving corresponding bounds.
Research on the cost of collisions is still in its early stages. Anderton et al. [1] show that collisions create substantial discrepancies between theoretical and observed performance in backoff algorithms, underscoring the need to account for collision costs. Building on this, Biswas et al. [4] provide one of the first analyses, addressing the problem of contention resolution and deriving bounds on expected latency and total expected collision cost.
1
2
3// Initialization
4 $w_{0}\leftarrow 2^{\mathcal{C}^{\epsilon}}$
5 $j\leftarrow 1$
6
7
8// Iterations
9 while true do
10
11 // Halving Phase
12 $w_{\mbox{\tiny cur}}\leftarrow w_{0}$
13 while $w_{\mbox{\tiny cur}}\geq 2$ do
14    for slot $i=1$ to $d\sqrt{\mathcal{C}}\,\ln(w_{\mbox{\tiny cur}})$ do
15
16
17     -2ptSend packet with probability $1/w_{cur}$
18    if success then
19        Terminate
20    else
21       $w_{\mbox{\tiny cur}}\leftarrow w_{\mbox{\tiny cur}}/2$
22
23
24 // Doubling Phase
25 $w_{\mbox{\tiny cur}}\leftarrow w_{0}$
26 for $s=1$ to $2^{j}$ do
27    $w_{\mbox{\tiny cur}}\leftarrow 2w_{\mbox{\tiny cur}}$
28    for slot $i=1$ to $d\sqrt{\mathcal{C}}\,\ln(w_{\mbox{\tiny cur}})$ do
29
30
31     -2ptSend packet with probability $1/w_{cur}$
32    if success then
33        Terminate
34
35 $j\leftarrow j+1$
Algorithm 1 Aim-High
3 Algorithm Specification
The pseudocode for AH is given above, and each packet executes this algorithm upon being activated. Under AH, each packet begins with an initial window size of $w_{0}$ $2^{\mathcal{C}^{\epsilon}}$.
[…]
After finishing a sample, if no success has occurred, then each packet sets $w_{\mbox{\tiny cur}}\leftarrow w_{\mbox{\tiny cur}}/2$. This is repeated until either a success (and termination) occurs, or further halving would reduce the window size below $2$ (Line 1). At which point, the doubling phase begins.
Doubling phase. At the beginning of this phase, $w_{\mbox{\tiny cur}}$ starts at $2w_{0}$. Each packet sends with probability $\Theta(1/w_{\mbox{\tiny cur}})$ in each slot of a sample (Line 1 and 1). If a success occurs, the execution terminates; otherwise, the window doubles and another sample is executed. This process continues until $2^{j}$ samples (each corresponding to a doubling of the window) are performed, where $j=1$ initially, and increments at the end of each iteration.
3.1 Context and Overview
We start by considering how prior approaches fare in a model when there is a collision cost. Then, we discuss our design choices regarding AH, along with an overview of our analysis.
3.1.1 A Baseline via Backoff.
A natural approach for breaking symmetry is to “backoff”, where each packet sends with probability $1/2^{i}$ in slot $i\geq 0$. At $i=\lg(n)$, the probability of a success in a slot becomes constant. An example of its use in solving wakeup is given by Jurdziński and Stachowiak [14], who achieve $O(n/\log n)$ latency with bounded error $\varepsilon$ in a similar model to ours (without a notion of collision cost); however, each active packet starts with a high (i.e., constant) sending probability, which inflicts $\Omega(\mathcal{C})$ cost.
So, we can observe that while backoff may be an important part of any solution, a key obstacle is how to set the initial sending probability suitably low. Since $n$ is unknown a priori, it seems that no matter what initial sending probability we choose—even if it depends on the known value of $\mathcal{C}$—the adversary may choose a sufficiently large value of $n$ such that collisions are unavoidable. What can we do?
3.1.2 Design Choices and Overview.
We now discuss how the design of AH allows us to do better than prior results in a model with collision cost.
Setting an Initial Window Size. The initial sending probability needs to be set carefully. Too high and we incur many collisions as we discussed above.
Consider a window of size $w^{*}=\Theta(n\sqrt{\mathcal{C}})$; we refer to this as the size of a good window. Informally, once packets reach a good window, the chance of a success in any slot of the window is at least $\binom{n}{1}(1/w^{*})\approx 1/\sqrt{\mathcal{C}}$. Therefore, over $\Theta(\sqrt{\mathcal{C}})$ slots, we expect a success. At the same time, the probability of a collision is approximately $\binom{n}{2}(1/w^{*})^{2}\approx 1/\mathcal{C}$, and thus the expected collision cost per slot is $1$, and $O(\sqrt{\mathcal{C}})$ over $\Theta(\sqrt{\mathcal{C}})$ slots. This observation dictates the size of our samples, which are $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ and appear in lines 1 and 1 in the halving phase, and lines 1 and 1 in the doubling phase of our pseudocode.
In designing AH, we hope that $w^{*}\leq w_{0}$, since in this case, the above reasoning yields good expected collision cost (and, as we will show, good latency). However, since the adversary sets $\mathcal{C}$ with knowledge of $n$, our hopes may be dashed; the adversary can always make collisions overwhelmingly likely. However, in this case, our consolation prize is that collisions are cheap, and the degree of cheapness corresponds to $w_{0}$. We choose to "aim high" with the initial window size; that is, setting $w_{0}=2^{\mathcal{C}^{\epsilon}}$. This choice sets up our analysis of the following two cases:
  1. 1.
    If $n\sqrt{\mathcal{C}}\leq w_{0}$, then the good-window size is less than or equal to the initial-window size. Here, the probability of a collision is not too high. Thus, so long as we adjust the sending probability upwards in a conservative manner, we can reduce expected collision cost; this is the purpose of the halving phase. Of course, there is a tension with latency, since this adjustment cannot be too conservative, otherwise the latency will be too large.
  2. 2.
    Else $n\sqrt{\mathcal{C}}>w_{0}$, and the good-window size is larger than the initial-window size. Here, the collision probability may be “too high”. However, a key realization is that $\mathcal{C}$ is small; we will argue that $n>2^{\mathcal{C}^{\epsilon}}/\sqrt{\mathcal{C}}$ implies that $\mathcal{C}=O(\log^{1/\epsilon}n)$. Thus, we can tolerate these collisions as we increase the window size in the doubling phase.
Why Iterations? Although AH is likely to terminate when a good window is reached, there is some small probability of failure. Suppose we are in Case 1, but this failure occurs in the halving phase. Then, this small probability of failure yields good expected total cost of collisions as the window size decreases past the good window down to $2$ (see Lemma 3). However, if the halving phase is never revisited, the packets may accrue intolerable latency by executing the doubling phase without success (as the sending probability drops off exponentially fast). Conversely, suppose we are in Case 2 and this failure occurs. Then, if we never reset $w_{\mbox{\tiny cur}}$ to (roughly) $w_{0}$, the latency costs may again be high. Therefore, iterations facilitate good expected latency (see Lemma 6).
Parameter $\epsilon$. Throughout, we treat $\epsilon$ as a small constant for the purposes of our analysis. Yet, it is worth pointing out how it interacts with the latency and the expected total cost of collisions. The former is improved by making $\epsilon$ small, since this shortens the halving phase. So, if aiming high pays off, then a smaller $\epsilon$ can reduce latency. However, we pay for this in terms of the latter metric; specifically, when $n\sqrt{\mathcal{C}}>w_{0}$ and the algorithm looks to succeed in the doubling phase, we have $\mathcal{C}=O(\log^{1/\epsilon}n)$. Thus, degree of the polynomial increases with the reciprocal of $\epsilon$.
4 Upper Bound
In this section, we analyze the performance of AH. Define $\eta$ $\max\{n,\mathcal{C}\}$. A good window was loosely discussed in Section 3.1.2; we define it as any window with size in $n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$. Throughout, we assume that $n$ is sufficiently large. We make use of the following well-known facts.
Fact 1
The following inequalities hold.
  1. (a)
    For any $x$, $1-x\leq e^{-x}$,
  2. (b)
    For any $0\leq x\leq 1/2$, $1-x\geq e^{-2x}$.
We also make use of the following standard Chernoff bounds.
Theorem 4.1
(Goemans [12]) Let $X=\sum_{i}X_{i}$, where $X_{i}$ are Bernoulli indicator random variables with probability $p$. The following holds:
$Pr\left(X\geq(1+\delta)E[X]\right)$ $\leq\exp\left(-\frac{\delta^{2}E[X]}{2+\delta}\right)\mbox{~~~for any $\delta>0$}$
and
[…]
The next lemma argues that success is likely in a good window.
Lemma 1
W.h.p. in $\eta$, a success occurs if AH executes in a good window.
Proof
Let $w^{*}\in[n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$ (recall Section 3). The probability of a success in any fixed slot of this good window is at least:
[…]
$\geq\left(\frac{1}{3e\sqrt{\mathcal{C}}}\right)\mbox{~~~since $\mathcal{C}\geq 4$.}$
Let $X_{i}=1$ if the $i$-th slot of the sample contains a success; otherwise, $X_{i}=0$. The expected number of successes in the sample of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots is at least:
$E[X]\geq\left(\frac{d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})}{3e\sqrt{\mathcal{C}}}\right)$ $=\left(\frac{d\ln(w_{\mbox{\tiny cur}})}{3e}\right)$
where $d^{\prime}$ is a constant that can be made arbitrarily large depending on $d$.
We then employ Theorem 4.1, with $\delta=1/2$, to obtain:
$Pr\left(X\leq(1-\delta)\frac{d\ln(w_{\mbox{\tiny cur}})}{3e}\right)$ $\leq Pr\left(X\leq\frac{d\ln(w_{\mbox{\tiny cur}})}{6e}\right)$
$\leq e^{-\frac{{(1/2)}^{2}d\ln(w_{\mbox{\tiny cur}})}{(5/2)3e}}$
$=e^{-\frac{d\ln(n\sqrt{\mathcal{C}})}{30e}}$
$\leq(n\sqrt{\mathcal{C}})^{-d^{\prime}}$
$\leq\eta^{-d^{\prime\prime}}$
where the third line follows from $w_{\mbox{\tiny cur}}=\Theta(\ln n\sqrt{\mathcal{C}})$, and $d^{\prime}$ and $d^{\prime\prime}$ can be arbitrarily large constants for $d$ is sufficiently large.
4.1 Case 1: $n\sqrt{\mathcal{C}}\leq w_{0}$.
In this case, the good window size is less than or equal to $w_{0}$, so we argue that AH is likely to obtain a success during the halving phase. We make use of the following known result that upper bounds the probability of a collision.
Lemma 2
(Biswas et al. [4]) Consider any slot $t$ where there exist $m$ active packets, each sending with the same probability $p$, where $p<1/m$. The probability of a collision in slot $t$ is at most $\frac{2m^{2}p^{2}}{\left(1-mp\right)}$.
Lemma 3
Assume that $n\sqrt{\mathcal{C}}\leq w_{0}$. The expected total cost of collisions for AH is $O(\mathcal{C}^{(1/2)+\epsilon})$.
Proof
[…]
$G+\sum_{j=1}^{\infty}q^{j}\cdot B_{j}$ (1)
Upper Bounding $B_{j}$. Clearly, $B_{j}$ is upper-bounded by the cost of having a collision in every single slot of iterations $1$ through $j$. In any iteration $k$, there is a sample of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots for each of the at most $\ln(w_{0})$ windows; thus, over iterations $1$ through $j$, the halving phases contributes $O(j\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\lg(w_{0}))$ slots. Additionally, in iteration $k$, the corresponding doubling phase executes $2^{k}$ samples, each consisting of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots; thus, by the sum of a geometric series, over iterations $1$ through $j$, the doubling phases contributing $O(\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})2^{j})$ slots. By the above:
$B_{j}$ $=O\left(j\,d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\lg(w_{0})+\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})2^{j}\right)\cdot\mathcal{C}$
$=O\left(j\,\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\log(w_{0})2^{j}\right)\cdot\mathcal{C}$
$=O(C^{(3/2)+2\epsilon}\,j^{2}\,2^{j})$
since $\log(w_{0})=\log(2^{\mathcal{C}^{\epsilon}})=O(\mathcal{C}^{\epsilon})$, and $\ln(w_{\mbox{\tiny cur}})\leq\ln(2^{j}w_{0})=O(j\mathcal{C}^{\epsilon})$.
By Lemma 1, the $q=1/\eta^{c}$, where recall that $\eta=\max\{n,\mathcal{C}\}$ where we can make the constant $c>0$ as large as we wish. Thus,
$\sum_{j=1}^{\infty}q^{j}\cdot B_{j}$ $\leq\sum_{j=1}^{\infty}(1/\eta^{c})^{j}\cdot O(C^{(3/2)+2\epsilon}\,j^{2}\,2^{j})$
$=O(1)$
where the last line follows for sufficiently large $c$.
Upper Bounding $G$. To upper bound $G$, for each slot $i$ in the first halving phase of the first iteration, let $X_{i}$ be an independent random variable where $X_{i}=1$ if the $i$-th slot contains a collision; otherwise, $X_{i}=0$. Recall that the windows of this first halving phase are repeatedly halved. For a slot in a window of size $w=w_{0}/2^{k}$, for a positive integer $k\geq 0$, by Lemma 2, with $m=n$ and $p=1/w$ (since $w\geq n\sqrt{\mathcal{C}}$ and $\mathcal{C}\geq 4$, we satisfy the requirement that $p<1/m$), we have:
$Pr(X_{i}=1)$ $\leq\frac{2(n/w)^{2}}{\left(1-(n/w)\right)}$
$=\frac{2(n/(w_{0}/2^{k}))^{2}}{\left(1-(n/(w_{0}/2^{k}))\right)}$
The expected cost due to collisions that occur in this particular window is thus at most:
$Pr(X_{i}=1)\cdot d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\cdot\mathcal{C}$ $\leq\frac{2(n/(w_{0}/2^{k}))^{2}}{\left(1-(n/(w_{0}/2^{k}))\right)}d\mathcal{C}^{(3/2)+\epsilon}.$
Note that the number of windows executed starting from an initial size $w_{0}=2^{\mathcal{C}^{\epsilon}}$ down to size $n\sqrt{\mathcal{C}}$, is at most $C^{\epsilon}-\lg{n}-(\lg\mathcal{C})/2$; for ease of presentation, let $x=C^{\epsilon}-\lg{n}-(\lg\mathcal{C})/2$. By the above, we can derive:
$G$ $\leq\sum_{k=0}^{x}\hskip 0.0pt\frac{2\left(\frac{n}{w_{0}/2^{k}}\right)^{2}}{\left(1-\left(\frac{n}{w_{0}/2^{k}}\right)\right)}\left(d\mathcal{C}^{(3/2)+\epsilon}\right)$
$=\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\sum_{k=0}^{x}\hskip 0.0pt\frac{2n^{2}\left(\frac{2^{k}}{2^{\mathcal{C}^{\epsilon}}}\right)^{2}}{\left(1-\frac{n2^{k}}{2^{\mathcal{C}^{\epsilon}}}\right)}$
$\leq\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{2n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{k=0}^{x}\hskip 0.0pt\frac{2^{2k}}{\left(1-\frac{1}{\sqrt{{\mathcal{C}}}}\right)}$ (2)
$\leq\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{k=0}^{x}\hskip 0.0pt2^{2k}$ (3)
$=\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\frac{4^{\mathcal{C}^{\epsilon}+1}}{(4-1)4^{\lg{n}}4^{\frac{\lg{\mathcal{C}}}{2}}}$ (4)
$<\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{8}{\mathcal{C}}$
$=O\left(\mathcal{C}^{(1/2)+\epsilon}\right).$
In the above, line 2 holds since $n\sqrt{\mathcal{C}}\leq 2^{\mathcal{C}^{\epsilon}}$, since we are in Case 1. Line 3 holds for any $\mathcal{C}\geq 4$, and line 4 holds by the sum of the geometric series. Thus, Equation 1 is bounded as claimed.
The next lemma bounds the expected latency of our algorithm, and its argument is similar to that of Lemma 3, although an extra $O(\mathcal{C}^{\epsilon})$-factor appears. Informally, this occurs since, in counting slots for the first halving phase, we have a sample size of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})=d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ $\leq d\sqrt{\mathcal{C}}\ln(w_{0})$ $d\sqrt{\mathcal{C}}C^{\epsilon}=O(\mathcal{C}^{(1/2)+\epsilon})$, and we have $\log_{2}w_{0}=O(C^{\epsilon})$ samples.
Lemma 4
Assume that $n\sqrt{\mathcal{C}}\leq w_{0}$. The expected latency of AH is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
4.2 Case 2: $n\sqrt{\mathcal{C}}>w_{0}$.
In this case, the good window size exceeds $w_{0}$, and so we argue that AH likely succeeds during the doubling phase. A key insight is that, for Case 2, $\mathcal{C}$ is small in terms of $n$.
Lemma 5
Assume $n\sqrt{\mathcal{C}}>w_{0}$. Then, $\mathcal{C}=O(\log^{1/\epsilon}n)$.
Proof
Since $n>2^{\mathcal{C}^{\epsilon}}/\sqrt{\mathcal{C}}$, we have:
$\lg n$ $>\mathcal{C}^{\epsilon}-(\lg\mathcal{C})/2$
$>\mathcal{C}^{\epsilon}-(\mathcal{C}^{\epsilon}/2).$ (5)
To justify the second line, let $\delta\in(0,1)$, and note that:
$(\lg\mathcal{C})/2$ $<\mathcal{C}^{\epsilon}/2$
so long as $\epsilon>\frac{\lg\lg(\mathcal{C})}{\lg\mathcal{C}}$, which holds by assumption since $\epsilon$ is a constant, and this is well-defined since $\mathcal{C}\geq 4$. Returning to Equation 5, we have that:
$(2\lg n)^{1/\epsilon}$ $>\mathcal{C}.$
as desired, since $\epsilon$ is a constant.
The next lemma establishes bounds on expected latency and total collision cost, and the intuition is as follows. In Case 2, collisions are “cheap” by Lemma 5, in the sense that $\mathcal{C}=O(\log^{1/\epsilon}n)$. Thus, the cost of collisions incurred until the good window is reached is small; it takes only $\log_{2}(n\sqrt{\mathcal{C}})+O(1)=O(\log n)$ doublings of the window size to reach a good window in the doubling phase. Since $2^{j}$ counts the number of doublings we perform in the doubling phase (recall Section 3), we reach a good window by $j=O(\log\log n)$. After this point the probability of success is high, and so subsequent iterations contribute little to the expected total cost of collisions.
Lemma 6
Assume $n\sqrt{\mathcal{C}}>w_{0}$. Then, the expected latency is $O\left(\log^{1/(2\epsilon)+4}n)\right)$.
Proof
The number of slots incurred for each individual halving phase is the sample size times the number of halvings, so upper bounded by:
$O\left(\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\right)\lg(w_{0})$ $=O\left(\sqrt{\mathcal{C}}\log(2^{j}w_{0})\log(w_{0})\right)$
$=O\left(\sqrt{\mathcal{C}}j\mathcal{C}^{2\epsilon}\right)$
$=O\left(j\log^{1/(2\epsilon)+2}n\right)$
where the last line follows by applying Lemma 5. Thus, over iterations $1$ to $j$, the number of slots contributed by the halving phases is $O(j^{2}\log^{1/(2\epsilon)+2}n)$
The number of slots in the doubling phase of the $j$-th iteration is:
$2^{j}d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ $=O(2^{j}\sqrt{\mathcal{C}}\log(2^{j}w_{0}))$
$=O(2^{j}\sqrt{\mathcal{C}}\,j\mathcal{C}^{\epsilon})$
$=O(2^{j}\,j\,(\log^{1/(2\epsilon)+1}n))$
slots, where the last line follow by Lemma 5. By the above, summing up the number of slots executed over iterations 1 to $j$ is $O(2^{j}\,j^{2}\,(\log^{1/(2\epsilon)+2}n))$.
Pessimistically, let us assume no success ever occurs during the halving phase. Focusing on the doubling phase, a good window is reached within at most $\log_{2}(n\sqrt{\mathcal{C}})+O(1)=\log_{2}n+O(\log\log n)$ doublings of the window size. Since $2^{j}$ counts the number of doublings we perform in the doubling phase (recall Section 3), we reach a good window by iteration $j=c^{\prime}\log\log n$, where $c^{\prime}>0$ is a sufficiently large constant. By the above, the total number of slots up until a good window is reached is:
$O\left(2^{j}\,j^{2}\,(\log^{1/(2\epsilon)+2}n)\right)$ $=O\left((\log n)\,(\log\log n)^{2}(\log^{1/(2\epsilon)+2}n)\right)$
$=O\left(\log^{1/(2\epsilon)+4}n\right)$
After this point the probability of success is high, by Lemma 1, and so subsequent iterations contribute little to the expected total cost of collisions. The expected latency is:
$O\left(\log^{1/(2\epsilon)+4}n\right)+\hskip-8.0pt\sum_{j=c^{\prime}\lg\lg n}^{\infty}\hskip-8.0ptO\left(\frac{1}{\eta}\right)^{j-c^{\prime}\lg\lg(n)+1}\hskip-5.0ptO\left(2^{j}\,j\,(\log^{1/(2\epsilon)+1}n)\right)$
$=O\left(\log^{1/(2\epsilon)+4}n\right)+\hskip-8.0pt\sum_{j=c^{\prime}\lg\lg n}^{\infty}\hskip-8.0ptO\left(\frac{1}{n}\right)^{j-c^{\prime}\lg\lg(n)+1}\hskip-10.0ptO\left(2^{j}\,j\,(\log^{1/(2\epsilon)+1}n)\right)$
$=O\left(\log^{1/(2\epsilon)+4}n\right)$
which completes the argument.
The following lemma about the expected total cost of collisions is proved similarly to Lemma 6. The main difference is that, even if every slot contains a collision, the cost increases only by a $\mathcal{C}=O(\log^{1/\epsilon}n)$ factor.
Lemma 7
Assume $n\sqrt{\mathcal{C}}>w_{0}$. Then, the expected total cost of collisions is $O\left(\log^{3/(2\epsilon)+4}n\right)$.
4.3 Performance of AH
We can now prove our upper bounds on the latency and expected collision cost for AH.
Theorem 1.1. AH solves the wakeup problem with the following costs:
  1. 1.
    If  $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, then the expected latency is $O(\mathcal{C}^{\frac{1}{2}+2\epsilon})$ and the expected total cost of collisions is $O(\mathcal{C}^{\frac{1}{2}+\epsilon})$.
  2. 2.
    Else, the expected latency is $O(\log^{4+\frac{1}{2\epsilon}}n)$ and the expected total cost of collisions is $O(\log^{4+\frac{3}{2\epsilon})}n)$.
Proof
The relationship between $n$ and $\mathcal{C}$ is established by the adversary (recall Section 1.1), and there are there are two cases to consider. In Case 1, $n\sqrt{\mathcal{C}}\leq w_{0}$. By Lemmas 3 and 4, the expected latency and the expected total collision cost for AH are each $O(\mathcal{C}^{(1/2)+2\epsilon})$, yielding (1).
In the second case, $n\sqrt{\mathcal{C}}>w_{0}$. By Lemmas, 6 and 7, the expected latency and expected total cost of collisions are $O(\log^{4+\frac{1}{2\epsilon}}n)$ and $O(\log^{4+\frac{3}{2\epsilon})}n)$, respectively, yielding (2).
5 Lower Bound
[…]
Lemma 8
Any fair algorithm for the wakeup problem that succeeds w.h.p. in $\eta$ must have total contention at least $2$.
Proof
[…]
$TC(\mathcal{A})=\sum_{u}LC(u)=n\cdot LC(u)$
By our assumption, $n\cdot LC(u)<2$. The probability of $u$ not succeeding in slot any fixed $t$ is at least 1p(t)1-p(t). Therefore, the probability that $u$ never succeeds over the execution of $\mathcal{A}$ is:
$\geq\prod_{i=1}^{\lambda}(1-p(i))$
$\geq e^{-2\sum_{i}^{\lambda}p(i)}$
$\geq e^{-2\cdot LC(u)}$
$\geq e^{-2\cdot TC(\mathcal{A})/n}$
$>e^{-4/n}$
where the second line follows from Fact 1(b), and the third line follows from the definition of $LC(u)$. The last line follows from the assumption that $TC(\mathcal{A})<2$.
It follows that the probability of all packets failing over the execution of $\mathcal{A}$ is:
$\geq\left(e^{-4/n}\right)^{n}=\frac{1}{e^{4}}$
Thus, there is a constant probability of all packets failing, which contradicts the w.h.p. guarantee of $\mathcal{A}$.
[…]
Lemma 9(Biswas et al. [5], Lemma 24)
Fix any slot $t$ in any fair algorithm for the wakeup problem. If $\texttt{Con}(t)>2$, then the probability of a collision is at least $1/10$.
By Lemma 9, any algorithm that has a slot with contention higher than $2$ has an expected collision cost that is $\Omega(\mathcal{C})$. Therefore, we need only to consider algorithms where every slot in the execution has contention at most $2$. Under these circumstances, we provide a lemma that lower bounds the probability of a collision as a function of contention. We use $\texttt{Con}^{2}(t)$ to denote $(\texttt{Con}(t))^{2}$.
Lemma 10
Fix any slot $t$ where there are at least $2$ packets and $\texttt{Con}(t)\leq 2$ under the execution of any fair algorithm for the wakeup problem. The probability of a collision in $t$ is $\Omega(\texttt{Con}^{2}(t))$.
Proof
For a fair algorithm, in any slot $t$, $\texttt{Con}(t)=np$. For any fixed slot, the probability of a collision involving exactly two packets (since there are at least $2$ packets by assumption) is a lower bound on the probability of a collision (that may involve more than $2$ packets), and this is $\binom{n}{2}p^{2}(1-p)^{n-2}$, which is at least:
$\geq\binom{n}{2}p^{2}(1-p)^{n-2}$
$\geq\binom{n}{2}p^{2}(1-p)^{n}\geq\left(\frac{n^{2}}{2}\right)\left(\frac{\texttt{Con}(t)}{n}\right)^{2}e^{-2\cdot\texttt{Con}(t)}\geq\left(\frac{\texttt{Con}^{2}(t)}{2e^{2\cdot\texttt{Con}(t)}}\right)$
$=\Omega(\texttt{Con}^{2}(t))$
where the last line follows from $\texttt{Con}(t)\leq 2$.
We can derive a relationship between latency and expected collision cost for fair algorithms.
Theorem 1.2. Let $\alpha$ and $\beta$ be constants such that $1/(2\epsilon)\geq\alpha\geq 0$ and $\beta\geq 0$. Any fair algorithm for the wakeup problem that has $O(\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n)$ latency has an expected total cost of collisions $\Omega(\mathcal{C}^{(1/2)-\alpha\epsilon}/\log^{\beta}n)$.
Proof
Let $\mathcal{A}$ be an algorithm that w.h.p. in $\eta$ solves wakeup with $L\leq d^{\prime}\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n$ latency, where $d^{\prime}>0$ is any constant. By Lemma 8, $\mathcal{A}$ has total contention:
$\sum_{t=1}^{L}\texttt{Con}(t)\geq 2$ (6)
As we discussed, an implication of Lemma 9, we need only consider that $\mathcal{A}$ has contention at most $2$ in every slot. Thus, we can use Lemma 10 (noting that $n\geq 2$), which implies that the expected cost of collisions in a single slot $t$ is:
[…]
for some constant $k>0$. So, the expected total cost of collisions is:
$\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)\mathcal{C}$
By Jensen’s inequality for convex functions, we have:
$\frac{\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)}{L}\geq\left(\frac{\sum_{t=1}^{L}k\,\texttt{Con}(t)}{L}\right)^{2}$
which implies:
$\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)\geq\frac{\left(\sum_{t=1}^{L}k\,\texttt{Con}(t)\right)^{2}}{L}$ (7)
So, by Equation 7, the expected cost is at least:
[…]
and since $L\leq d^{\prime}\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n$, we have:
$\frac{\mathcal{C}\,k^{2}\,2^{2}}{L}$ $=\Omega(\mathcal{C}^{(1/2)-\alpha\epsilon}/\log^{\beta}n)$
which yields the claim.
As evidence that our result is not far off from the asymptotic optimal, we note that any fair algorithms with latency $O(\mathcal{C}^{\frac{1}{2}+2\epsilon})$ must have an expected total cost of collisions that is $\Omega(\mathcal{C}^{\frac{1}{2}-2\epsilon})$, which is an $O(\mathcal{C}^{3\epsilon})$-factor smaller than that achieved by AH.
6 Conclusion and Future Work
The theoretical exploration of collision costs remains in its early stages. Our work here provides steps towards addressing fundamental problems that connect with the wireless domain.
For future work, we aim to prove more general lower bounds. Additionally, we wish to derive results for the dynamic setting, where packets arrive over time rather than all being initially present.